# Quaternion-Domain Super MDS for Robust 3D Localization

### v1

@@ Abstract
This paper proposes a novel low-complexity three-dimensional (3D) localization algorithm for wireless sensor networks, termed quanternion-domain super multi-dimensional scaling (QD-SMDS). The algorithm is based on a reformulation of the SMDS, originally developed in the real domain, using quaternion algebra. By representing 3D coordinates as quaternions, the method constructs a rank-$1$ Gram edge kernel (GEK) matrix that integrates both relative distance and angular information between nodes, which enhances the noise reduction effect achieved through low-rank truncation employing singular value decomposition (SVD), thereby improving robustness against information loss. To further reduce computational complexity, we also propose a variant of QD-SMDS that eliminates the need for the computationally expensive SVD by leveraging the inherent structure of the quaternion-domain GEK matrix. This alternative directly estimates node coordinates using only matrix multiplications within the quaternion domain. Simulation results demonstrate that the proposed method significantly improves localization accuracy compared to the original SMDS algorithm, especially in scenarios with substantial measurement errors. The proposed method also achieves comparable localization accuracy without requiring SVD.
@@ I INTRODUCTION
With recent advancements in sensor technology, wireless sensor networks (WSNs) have increasingly emerged as fundamental information infrastructures across various industrial domains [1, 2, 3], including precision agriculture [4], smart factories [5], and medical sensing [6]. Among different types of sensor data, location information plays a particularly vital role—not only enhancing the value of the sensed data itself, but also contributing significantly to the efficient operation of WSNs. In fact, its importance is often regarded as comparable to that of payload data in conventional wireless communications. Many services utilizing such location information are designed under the assumption of large-scale networks consisting of numerous small, low-power sensor terminals (referred to as nodes[7, 8]. In these scenarios, there is a strong demand for algorithms capable of simultaneously estimating the positions of multiple nodes with high accuracy and low computational complexity, based on the aggregated multidimensional information available from WSNs [9, 10, 11].
@@
Localization algorithms can be broadly categorized into three main approaches based on their underlying mathematical frameworks: Bayesian inference [12], convex optimization [13], and isometric embedding methods. When selecting an appropriate algorithm, it is crucial to consider the trade-off between computational complexity and estimation accuracy. This paper focuses on the isometric embedding approach, commonly known as multidimensional scaling (MDS) [14]. Unlike Bayesian inference and convex optimization, MDS offers the notable advantage of fixed computational complexity. While the convergence speed of Bayesian and optimization-based methods can vary significantly due to measurement noise or missing data, MDS maintains consistent complexity regardless of such factors. This makes MDS particularly well-suited for systems with limited computational resources and strict time constraints.
@@
Among various MDS-based localization methods, super multi-dimensional scaling (SMDS) [16, 17] stands out as a hybrid algorithm that integrates both distance and angle information. It has demonstrated superior performance over classical MDS —which relies solely on distance information—even in the presence of angular uncertainties of approximately ±$\pm 35^{\circ}$  [18]. Complex-domain SMDS (CD-SMDS) is a technique that reduces computational complexity and improves precision by tailoring SMDS for two-dimensional (2D) localization [19]. Unlike conventional SMDS, which represents 2D node coordinates as real-valued vectors, CD-SMDS uses complex-valued scalars to express node positions. By constructing a GEK matrix in the complex domain that consolidates all measurement data, the method enables rank reduction to one and enhances accuracy through noise suppression via low-rank truncation using SVD [20]. This demonstrates that translating real-valued vectors into single complex scalars not only simplifies representation but also provides a foundation for extending SMDS to a 3D localization framework. In other words, if 3D node coordinates can similarly be expressed as scalars, it may be possible to develop an SMDS algorithm optimized for 3D localization.
@@
Quaternions are an extension of complex numbers and are widely used to represent 3D rotations in fields such as computer graphics and computer vision  [27, 26]. A quaternion consists of one real component and three imaginary components  [22]. By utilizing three of its four degrees of freedom (DoF), it becomes possible to represent 3D node coordinates and construct a rank-$1$ GEK matrix in the quaternion domain. The algebraic structure of quaternions is well established, and frameworks for performing algebraic operations such as SVD are also available [23]. This enables us to exploit SVD-based low-rank truncation for noise suppression and thereby improve localization accuracy—leading to the development of QD-SMDS.
However, the computational efficiency gained by QD-SMDS is quite limited and may even result in increased computational cost in some systems. This is because performing SVD in the quaternion domain requires converting the target quaternion matrix into a complex-valued matrix of double the size, followed by applying SVD to the resulting complex-equivalent matrix. Given that the size of the GEK matrix increases proportionally with the number of node pairs, this process can become impractical—particularly in systems with constrained computational resources. To eliminate the need for SVD in SMDS, an iterative approach based on maximal ratio combining (MRC) was proposed in [19]. However, due to the non-commutative nature of quaternion multiplication, directly extending this method to QD-SMDS is not straightforward. In this paper, we address this challenge by explicitly considering quaternion non-commutativity in the combining operation and propose a novel, low-complexity 3D localization algorithm—termed quaternion-domain maximal ratio combining super multidimensional scaling (QD-MRC-SMDS) —which is constructed solely from simple quaternion multiplications. The contributions of this paper are summarized as follows [^1] The conference paper [24] is an earlier version of this paper, which was presented at the IEEE SPAWC 2025.:
@@
[…]
@@ II-A Basics of Quaternion Algebra
[…]
@@ III-A Problem Formulation
where $\bm{C}\triangleq\left[\bm{C}_{\mathrm{AA}},\bm{C}_{\mathrm{AT}}\right]^{\mathsf{T}}\in\mathbb{R}^{M\times N}$ is a structure matrix encoding the pairwise relationships between nodes and edges.
@@ III-B SMDS Algorithm Recap
$\bm{K}_{\mathrm{r}}\!\!\!$$\triangleq$ $\!\!\!\bm{V}\bm{V}^{\mathsf{T}}$ (22)
$=$ $\!\!\!\mathrm{diag}\left(\bm{d}\right)\begin{bmatrix}\cos\alpha_{11}\!\!\!\!&\!\!\!\!\cdots\!\!\!\!&\!\!\!\!\cos\alpha_{1M}\\ \vdots&\ddots&\vdots\\ \cos\alpha_{M1}\!\!\!\!&\!\!\!\!\cdots\!\!\!\!&\!\!\!\!\cos\alpha_{MM}\end{bmatrix}\mathrm{diag}\left(\bm{d}\right),$
@@
Algorithm 1 SMDS Algorithm
1:Input:
3:The coordinates of at least $4$ ANs.
4:Steps:
6:Perform SVD of the constructed GEK matrix $\tilde{\bm{K}}_{\mathrm{r}}$.
[…]
8:Compute $\hat{\bm{X}}$ from $\hat{\bm{V}}$ in Eq. (24).
@@
Finally, the estimated coordinate matrix can be recovered from $\hat{\bm{V}}$ by inverting the relationship in (19), i.e., $\hat{\bm{X}}=\bm{C}^{-1}\hat{\bm{V}}$. However, since the rank of the structure matrix $\bm{C}$, as defined in (20), is $N-1$, this inverse cannot be directly computed. To circumvent this rank deficiency, we exploit the known coordinate matrix $\bm{X}_{\mathrm{A}}$ corresponding to ANs in (14). By incorporating this prior knowledge, the inversion can be reformulated as
$\left[\begin{array}[]{c}\bm{X}_{\mathrm{A}}\\ \hline\cr\hat{\bm{X}}\end{array}\right]=\left[\begin{array}[]{c|c}\bm{I}_{N_{\mathrm{A}}}&\bm{0}_{N_{\mathrm{A}}\times N_{\mathrm{T}}}\\ \hline\cr\bm{0}_{M\times N_{A}}&\bm{C}\end{array}\right]^{-1}\left[\begin{array}[]{c}\bm{X}_{\mathrm{A}}\\ \hline\cr\hat{\bm{V}}\end{array}\right].$ (24)
@@ IV-A1 Derivation of the Proposed Algorithm
where out of the four DoF, three are used for the $(x,y,z)$ coordinates, and the remaining one is set to 00222In this paper, the component corresponding to the imaginary unit $\mathbf{k}$ is set to zero. Although any one of the four components can be chosen as zero, it should be noted that the GEK matrix must be modified accordingly if a different component is selected. .
@@
$\bm{v}_{m},\bm{v}_{p}\right>\triangleq\left[\grave{a}_{m},\grave{b}_{m},\grave{c}_{m}\right]\begin{bmatrix}\grave{a}_{p}\\ \grave{b}_{p}\\ \grave{c}_{p}\end{bmatrix}$
@@
The corresponding Euclidean norms (distances) are give by $d_{m}^{(\mathrm{xy})}\triangleq\|\bm{v}_{m}^{(\mathrm{xy})}\|$, $d_{m}^{(\mathrm{xz})}\triangleq\|\bm{v}_{m}^{(\mathrm{xz})}\|$, and $d_{m}^{(\mathrm{yz})}\triangleq\|\bm{v}_{m}^{(\mathrm{yz})}\|$. Similarly, $\alpha_{mp}^{(\mathrm{xy})}$, $\alpha_{mp}^{(\mathrm{xz})}$, and $\alpha_{mp}^{(\mathrm{yz})}$ represent the ADoAs between two 2D projected vectors on the $(\mathrm{x},\mathrm{y})$, $(\mathrm{x},\mathrm{z})$, and $(\mathrm{y},\mathrm{z})$ planes, respectively. For clarity of these parameters, please refer to Fig.1 at the top of this page.
@@
$\nu_{m}\nu_{p}^{*}$ $=\underbrace{\left(\grave{a}_{m}\grave{a}_{p}+\grave{b}_{m}\grave{b}_{p}+\grave{c}_{m}\grave{c}_{p}\right)}_{\left<\bm{v}_{m},\bm{v}_{p}\right>}+\mathbf{i}\underbrace{\left(\grave{a}_{p}\grave{b}_{m}-\grave{a}_{m}\grave{b}_{p}\right)}_{-\left|\bm{v}_{m}^{(\mathrm{xy})}\times\bm{v}_{p}^{(\mathrm{xy})}\right|}$ (31)
@@
$\bm{K}_{\mathrm{q}}\triangleq\bm{\nu}\bm{\nu}^{\mathsf{H}}=\mathrm{diag}\left[\bm{d}\right]\begin{bmatrix}\cos\alpha_{11}&\cdots&\cos\alpha_{1M}\\ \vdots&\ddots&\vdots\\ \cos\alpha_{M1}&\cdots&\cos\alpha_{MM}\end{bmatrix}\mathrm{diag}\left[\bm{d}\right]-\mathbf{i}\;\mathrm{diag}\left[\bm{d}^{(\mathrm{xy})}\right]\begin{bmatrix}\sin\alpha_{11}^{(\mathrm{xy})}\!\!\!&\!\!\!\cdots\!\!\!&\!\!\!\sin\alpha_{1M}^{(\mathrm{xy})}\\ \vdots&\ddots&\vdots\\ \sin\alpha_{M1}^{(\mathrm{xy})}\!\!\!&\!\!\!\cdots\!\!\!&\!\!\!\sin\alpha_{MM}^{(\mathrm{xy})}\end{bmatrix}\mathrm{diag}\left[\bm{d}^{(\mathrm{xy})}\right]$
$\quad-\mathbf{j}\;\mathrm{diag}\left[\bm{d}^{(\mathrm{xz})}\right]\begin{bmatrix}\sin\alpha_{11}^{(\mathrm{xz})}\!\!\!&\!\!\!\cdots\!\!\!&\!\!\!\sin\alpha_{1M}^{(\mathrm{xz})}\\ \vdots&\ddots&\vdots\\ \sin\alpha_{M1}^{(\mathrm{xz})}\!\!\!&\!\!\!\cdots\!\!\!&\!\!\!\sin\alpha_{MM}^{(\mathrm{xz})}\end{bmatrix}\mathrm{diag}\left[\bm{d}^{(\mathrm{xz})}\right]-\mathbf{k}\;\mathrm{diag}\left[\bm{d}^{(\mathrm{yz})}\right]\begin{bmatrix}\sin\alpha_{11}^{(\mathrm{yz})}\!\!\!&\!\!\!\cdots\!\!\!&\!\!\!\sin\alpha_{1M}^{(\mathrm{yz})}\\ \vdots&\ddots&\vdots\\ \sin\alpha_{M1}^{(\mathrm{yz})}\!\!\!&\!\!\!\cdots\!\!\!&\!\!\!\sin\alpha_{MM}^{(\mathrm{yz})}\end{bmatrix}\mathrm{diag}\left[\bm{d}^{(\mathrm{yz})}\right].$ (32)
From (31), the rank-$1$ quaternion-domain GEK matrix, which incorporates all pairwise distances and ADoA information among the nodes, is given in (IV-A1), shown at the top of the next page. Assuming that measured values of all pairwise distance and ADoA parameters appearing in (31) are available, the quaternion-domain GEK matrix with measurement errors, denoted as $\tilde{\bm{K}}_{\mathrm{q}}$, can be constructed. Accordingly, the estimate of the quaternion edge vector $\bm{\nu}$ is given as
@@
The SVD of $\tilde{\bm{K}}_{\mathrm{q}}$ is performed using the QSVD introduced in Section II-B. Note that, in our case, since the quaternion-domain GEK $\tilde{\bm{K}}_{\mathrm{q}}$ is regular, the quaternion eigendecomposition is effectively equivalent to the QSVD.

### v2

@@ Abstract
This paper proposes a novel low-complexity three-dimensional (3D) localization algorithm for wireless sensor networks, termed quanternion-domain super multi-dimensional scaling (QD-SMDS). The algorithm is based on a reformulation of the SMDS, originally developed in the real domain and for two-dimensional (2D) problems, using quaternion algebra. By representing 3D coordinates as quaternions, the method constructs a rank-$1$ Gram edge kernel (GEK) matrix that integrates both relative distance and angular information between nodes, which enhances the noise reduction effect achieved through low-rank truncation employing singular value decomposition (SVD), thereby improving robustness against information loss. To further reduce computational complexity, we also propose a variant of QD-SMDS that eliminates the need for the computationally expensive SVD by leveraging the inherent structure of the quaternion-domain GEK matrix. This alternative directly estimates node coordinates using only matrix multiplications within the quaternion domain. Simulation results demonstrate that the proposed method significantly improves localization accuracy compared to the original SMDS algorithm, especially in scenarios with substantial measurement errors. The proposed method also achieves comparable localization accuracy without requiring SVD.
@@ I INTRODUCTION
With recent advancements in sensor technology, wireless sensor networks (WSNs) have increasingly emerged as fundamental information infrastructures across a wide range of industrial domains [1, 2, 3], including precision agriculture [4], smart factories [5], and medical sensing [6]. Among the various types of sensor data, location information plays a particularly vital role, as it not only enhances the value of the sensed data itself but also contributes significantly to the efficient operation of WSNs. In fact, its importance is often regarded as comparable to that of payload data in conventional wireless communications. Accordingly, many services that utilize such location information are designed under the assumption of large-scale networks composed of numerous small, low-power sensor terminals (hereafter referred to as nodes[7, 8].
In these scenarios, there is a strong demand for algorithms capable of simultaneously estimating the positions of multiple nodes with high accuracy and low computational complexity, based on the aggregated multidimensional information available from WSNs [9, 10, 11].
@@
Localization algorithms can be broadly categorized into three main approaches based on their underlying mathematical frameworks: Bayesian inference [12, 13, 14, 15], convex optimization [16], and isometric embedding methods. When selecting an appropriate algorithm, it is crucial to consider the trade-off between computational complexity and estimation accuracy. This paper focuses on the isometric embedding approach, commonly known as multidimensional scaling (MDS) [17]. Unlike Bayesian inference and convex optimization, MDS offers the notable advantage of fixed computational complexity. While the convergence speed of Bayesian and optimization-based methods can vary significantly due to measurement noise or missing data, MDS maintains consistent complexity regardless of such factors. This makes MDS particularly well-suited for systems with limited computational resources and strict time constraints.
@@
Among various MDS-based localization methods, super multi-dimensional scaling (SMDS) [19, 20] stands out as a hybrid algorithm that integrates both distance and angle information. It has demonstrated superior performance over classical MDS —which relies solely on distance information—even in the presence of angular uncertainties of approximately $\pm 35^{\circ}$  [21]. Complex-domain SMDS (CD-SMDS) is a technique that reduces computational complexity and improves precision by tailoring SMDS for two-dimensional (2D) localization [22]. Unlike conventional SMDS, which represents 2D node coordinates as real-valued vectors, CD-SMDS uses complex-valued scalars to express node positions. By constructing a Gram edge kernel (GEK) matrix in the complex domain that consolidates all measurement data, the method enables rank reduction to one and enhances accuracy through noise suppression via low-rank truncation using singular value decomposition (SVD) [23]. This demonstrates that translating real-valued vectors into single complex scalars not only simplifies representation but also provides a foundation for extending SMDS to a three-dimensional (3D) localization framework. In other words, if 3D node coordinates can be similarly represented as scalars, it may be possible to develop an SMDS algorithm optimized for 3D localization that reduces the rank of the GEK matrix to one, favoring matrix completion techniques, and thereby improves localization accuracy, as demonstrated by CD-SMDS for 2D coordinates.
@@
Quaternions constitute the smallest and most widely established extension of complex numbers that preserves an analogous algebraic structure and are widely used to represent 3D rotations in fields such as computer graphics and computer vision [30, 29]. A quaternion consists of one real component and three imaginary components [25]. By utilizing three of its four degrees of freedom (DoF), it becomes possible to represent 3D node coordinates and construct a rank-$1$ GEK matrix in the quaternion domain. Furthermore, a mature matrix toolbox for quaternions is available, providing well-defined conjugation and norm operations, as well as frameworks for performing algebraic computations such as SVD [26]. This enables the exploitation of SVD-based low-rank truncation for noise suppression, thereby improving localization accuracy and leading to the development of quanternion-domain super multi-dimensional scaling (QD-SMDS).
However, the computational efficiency gained by QD-SMDS is quite limited and may even result in increased computational cost in some systems. This is because performing SVD in the quaternion domain requires converting the target quaternion matrix into a complex-valued matrix of double the size, followed by applying SVD to the resulting complex-equivalent matrix. Given that the size of the GEK matrix increases proportionally with the number of node pairs, this process can become impractical—particularly in systems with constrained computational resources. To eliminate the need for SVD in SMDS, an iterative approach based on maximal ratio combining (MRC) was proposed in [22]. However, due to the non-commutative nature of quaternion multiplication, directly extending this method to QD-SMDS is not straightforward. In this paper, we address this challenge by explicitly considering quaternion non-commutativity in the combining operation and propose a novel, low-complexity 3D localization algorithm—termed quaternion-domain maximal ratio combining super multidimensional scaling (QD-MRC-SMDS) —which is constructed solely from simple quaternion multiplications. The contributions of this paper are summarized as follows [^1] The conference paper [27] is a shorter and earlier version of this work and was presented at the IEEE SPAWC 2025. In particular, [27] introduced the basic concept of QD-SMDS and and provided a brief performance evaluation by comparing it with the preceding SMDS. In contrast, the present paper offers a detailed derivation of QD-SMDS and introduces a new lower-complexity variant, QD-MRC-SMDS, together with a comprehensive performance assessment.:
@@
Notation: The following notation is used throughout unless otherwise specified. Sets of real and complex numbers are denoted by $\mathbb{R}$ and $\mathbb{C}$, respectively. Vectors and matrices are denoted by lower- and upper-case bold-face letters, respectively. The conjugate, transpose, and conjugate transpose operators are denoted by $(\cdot)^{*}$, $(\cdot)^{\mathsf{T}}$, and $(\cdot)^{\mathsf{H}}$, respectively. The $a\times a$ square identity matrix is denoted by $\bm{I}_{a}$. The $a\times b$ all-zeros matrix and all-ones matrix are denoted by $\bm{0}_{a\times b}$ and $\bm{1}_{a\times b}$, respectively. The diagonal matrix constructed by placing the elements of a vector $\bm{a}$ on its main diagonal is denoted by $\mathrm{diag}\left[\bm{a}\right]$. The Euclidean norm and Frobenius norm are denoted by $\|\cdot\|$ and $\|\cdot\|_{\mathrm{F}}$, respectively. The inner product and the outer product are denoted by $\left$$\cdot$$\cdot\right$> and $\left|\cdot\times\cdot\right|$, respectively. The determinant of a matrix $\bm{A}$ is denoted by $\mathrm{det}\left[\bm{A}\right]$. The matrix formed by stacking the odd-numbered rows of a matrix $\bm{A}$ is denoted by $\mathrm{r}_{\mathrm{odd}}\left[\bm{A}\right]$, and the matrix formed by collecting the odd-numbered columns is denoted by $\mathrm{c}_{\mathrm{odd}}\left[\bm{A}\right]$. The element-wise $k$-th power of the entries in a matrix $\bm{A}$ is denoted by $\bm{A}^{\odot k}$.
@@ II-A Basics of Quaternion Algebra
$\mathbf{i}\cdot\mathbf{j}\!\!$ $\!\!=\!\!$ $\!\!-\mathbf{j}\cdot\mathbf{i}=\mathbf{k},$ (3a)
$\mathbf{j}\cdot\mathbf{k}\!\!$ $\!\!=\!\!$ $\!\!-\mathbf{k}\cdot\mathbf{j}=\mathbf{i},$ (3b)
$\mathbf{k}\cdot\mathbf{i}\!\!$ $\!\!=\!\!$ $\!\!-\mathbf{i}\cdot\mathbf{k}=\mathbf{j}.$ (3c)
@@ III-A Problem Formulation
where $\bm{C}\triangleq\left[\bm{C}_{\mathrm{AA}}^{\mathsf{T}},\bm{C}_{\mathrm{AT}}^{\mathsf{T}}\right]^{\mathsf{T}}\in\mathbb{R}^{M\times N}$ is a structure matrix encoding the pairwise relationships between nodes and edges.
@@ III-B SMDS Algorithm Recap
$\hskip-20.00003pt\bm{K}_{\mathrm{r}}\!\!\!$ $\triangleq$ $\!\!\!\bm{V}\bm{V}^{\mathsf{T}}$ (22)
$=$ $\!\!\!\mathrm{diag}\left(\bm{d}\right)\begin{bmatrix}\cos\alpha_{11}\!\!\!\!&\!\!\!\!\cdots\!\!\!\!&\!\!\!\!\cos\alpha_{1M}\\[-4.30554pt] \vdots&\ddots&\vdots\\ \cos\alpha_{M1}\!\!\!\!&\!\!\!\!\cdots\!\!\!\!&\!\!\!\!\cos\alpha_{MM}\end{bmatrix}\mathrm{diag}\left(\bm{d}\right),$
@@
Finally, the estimated coordinate matrix can be recovered from $\hat{\bm{V}}$ by inverting the relationship in (19), i.e., $\hat{\bm{X}}=\bm{C}^{-1}\hat{\bm{V}}$. However, since the rank of the structure matrix $\bm{C}$, as defined in (20), is N1N-1, this inverse cannot be directly computed. To circumvent this rank deficiency, we exploit the known coordinate matrix $\bm{X}_{\mathrm{A}}$ corresponding to ANs in (14). By incorporating this prior knowledge, the inversion can be reformulated as
[…]
@@
The computational complexity of the SMDS algorithm is dominated by the truncated SVD of the kernel matrix in (22), in which only the three dominant eigenpairs are computed [32]. Accordingly, the asymptotic complexity of this approach is $O\left(3M^{2}\right)=O\left((N_{A}(N_{A}-1)/2+N_{A}N_{T})^{2}\right)=O\left(N_{A}^{2}(N_{A}+N_{T})^{2}\right)$.
Algorithm 1 SMDS Algorithm
1:Input:
3:The coordinates of at least 44 ANs.
4:Steps:
6:Perform SVD of the constructed GEK matrix $\tilde{\bm{K}}_{\mathrm{r}}$.
7:Obtain the edge matrix $\hat{\bm{V}}$ using Eq. (23)
8:Compute $\hat{\bm{X}}$ from $\hat{\bm{V}}$ in Eq. (24).
@@
III-C Classical Localization Models and Link to SMDS and its Quaternion Variants
In classical WSN localization, unknown node positions $\bm{X}_{\mathrm{T}}$ are estimated using range-based models, angle-based models (e.g., angle of arrival (AoA)/direction of arrival (DoA)), and hybrid models that fuse distance and angle information [8, 11, 13]. Both SMDS and in its quaternion-domain variants adopt such a hybrid framework, in which range estimates, ADoAs between edge vectors, and DoAs are jointly used to estimate node locations.
Given these measurements, most classical localization methods estimate $\bm{X}_{\mathrm{T}}$ by solving a nonlinear optimization problem, for example via maximum likelihood/least-squares formulations, Bayesian filtering or message passing techniques, or convex relaxation approaches such as semidefinite programming [11, 16, 46, 13]. In contrast, SMDS transforms the same measurement information into a structured matrix problem. Specifically, distance and angle measurements are used to construct a low-rank GEK matrix, from which edge-vector estimates and, subsequently, the target locations $\bm{X}_{\mathrm{T}}$ are recovered. This separation of steps (measurement acquisition $\rightarrow$ GEK construction $\rightarrow$ low-rank recovery $\rightarrow$ coordinate reconstruction) is a key feature that is exploited by the proposed quaternion-domain extensions.
The proposed quaternion-domain methods preserve the same sensing model and localization objective while modifying the algebraic representation of 3D edges from real vectors to quaternions. This modification alters the structure and rank of the corresponding GEK matrix and will be shown to yield more robust position estimates, particularly in the presence of strong measurement noise and missing information.
@@ IV-A1 Derivation of the Proposed Algorithm
where out of the four DoF, three are used for the $(x,y,z)$ coordinates, and the remaining one is set to 0222The choice of embedding the 3D coordinates into these components is arbitrary. In principle, any three of the four components can be used, and the resulting GEK matrix retains the same structure, with the angle and distance matrices appearing with different imaginary units. When extracting the estimated coordinates, the components corresponding to the three imaginary units selected to represent the $(x,y,z)$ coordinates must then be taken..
@@
$\!\!\!\!\left<\bm{v}_{m},\bm{v}_{p}\right>\triangleq\left[\grave{a}_{m},\grave{b}_{m},\grave{c}_{m}\right]\begin{bmatrix}\grave{a}_{p}\\ \grave{b}_{p}\\ \grave{c}_{p}\end{bmatrix}$
@@
For clarity, the notation for the real and quaternion vectors, together with their corresponding projections, is summarized in Table I.
TABLE I: Edge Vector Notation used in Real and Quaternion Domains.
Symbol Meaning / domain
$\bm{v}_{m}$ $m$-th edge vector in $\mathbb{R}^{3}$ (e.g., $\bm{v}_{m}=[\grave{a}_{m},\,\grave{b}_{m},\,\grave{c}_{m}]^{\mathsf{T}}$)
$\nu_{m}$ Quaternion representation of $\bm{v}_{m}$; $\nu_{m}=\grave{a}_{m}+\mathbf{i}\grave{b}_{m}+\mathbf{j}\grave{c}_{m}+\mathbf{k}\cdot 0$
$\bm{v}_{m}^{(xy)}$ xyxy-projection of $\bm{v}_{m}$ in $\mathbb{R}^{2}$ (e.g., $\bm{v}_{m}^{(xy)}=[\grave{a}_{m},\,\grave{b}_{m}]^{\mathsf{T}}$)
$\bm{v}_{m}^{(xz)}$ xzxz-projection of $\bm{v}_{m}$ in $\mathbb{R}^{2}$
$\bm{v}_{m}^{(yz)}$ yzyz-projection of $\bm{v}_{m}$ in $\mathbb{R}^{2}$
The corresponding Euclidean norms (distances) are given by $d_{m}^{(\mathrm{xy})}\triangleq\|\bm{v}_{m}^{(\mathrm{xy})}\|$, $d_{m}^{(\mathrm{xz})}\triangleq\|\bm{v}_{m}^{(\mathrm{xz})}\|$, and $d_{m}^{(\mathrm{yz})}\triangleq\|\bm{v}_{m}^{(\mathrm{yz})}\|$. Similarly, $\alpha_{mp}^{(\mathrm{xy})}$, $\alpha_{mp}^{(\mathrm{xz})}$, and $\alpha_{mp}^{(\mathrm{yz})}$ represent the ADoAs between two 2D projected vectors on the $(\mathrm{x},\mathrm{y})$, $(\mathrm{x},\mathrm{z})$, and $(\mathrm{y},\mathrm{z})$ planes, respectively. For clarity of these parameters, please refer to Fig.1 at the top of this page.
@@
$\hskip-21.52771pt\nu_{m}\nu_{p}^{*}\hskip-17.22217pt$ $=\underbrace{\left(\grave{a}_{m}\grave{a}_{p}+\grave{b}_{m}\grave{b}_{p}+\grave{c}_{m}\grave{c}_{p}\right)}_{\left<\bm{v}_{m},\bm{v}_{p}\right>}+\mathbf{i}\underbrace{\left(\grave{a}_{p}\grave{b}_{m}-\grave{a}_{m}\grave{b}_{p}\right)}_{-\left|\bm{v}_{m}^{(\mathrm{xy})}\times\bm{v}_{p}^{(\mathrm{xy})}\right|}$ (31)
@@
$\bm{K}_{\mathrm{q}}\triangleq\bm{\nu}\bm{\nu}^{\mathsf{H}}=\mathrm{diag}\left[\bm{d}\right]\begin{bmatrix}\cos\alpha_{11}&\cdots&\cos\alpha_{1M}\\ \vdots&\ddots&\vdots\\ \cos\alpha_{M1}&\cdots&\cos\alpha_{MM}\end{bmatrix}\mathrm{diag}\left[\bm{d}\right]-\mathbf{i}\;\mathrm{diag}\left[\bm{d}^{(\mathrm{xy})}\right]\begin{bmatrix}\sin\alpha_{11}^{(\mathrm{xy})}\!\!\!&\!\!\!\cdots\!\!\!&\!\!\!\sin\alpha_{1M}^{(\mathrm{xy})}\\ \vdots&\ddots&\vdots\\ \sin\alpha_{M1}^{(\mathrm{xy})}\!\!\!&\!\!\!\cdots\!\!\!&\!\!\!\sin\alpha_{MM}^{(\mathrm{xy})}\end{bmatrix}\mathrm{diag}\left[\bm{d}^{(\mathrm{xy})}\right]$ (32)
$\quad-\mathbf{j}\;\mathrm{diag}\left[\bm{d}^{(\mathrm{xz})}\right]\begin{bmatrix}\sin\alpha_{11}^{(\mathrm{xz})}\!\!\!&\!\!\!\cdots\!\!\!&\!\!\!\sin\alpha_{1M}^{(\mathrm{xz})}\\ \vdots&\ddots&\vdots\\ \sin\alpha_{M1}^{(\mathrm{xz})}\!\!\!&\!\!\!\cdots\!\!\!&\!\!\!\sin\alpha_{MM}^{(\mathrm{xz})}\end{bmatrix}\mathrm{diag}\left[\bm{d}^{(\mathrm{xz})}\right]-\mathbf{k}\;\mathrm{diag}\left[\bm{d}^{(\mathrm{yz})}\right]\begin{bmatrix}\sin\alpha_{11}^{(\mathrm{yz})}\!\!\!&\!\!\!\cdots\!\!\!&\!\!\!\sin\alpha_{1M}^{(\mathrm{yz})}\\ \vdots&\ddots&\vdots\\ \sin\alpha_{M1}^{(\mathrm{yz})}\!\!\!&\!\!\!\cdots\!\!\!&\!\!\!\sin\alpha_{MM}^{(\mathrm{yz})}\end{bmatrix}\mathrm{diag}\left[\bm{d}^{(\mathrm{yz})}\right].$
From (31), the rank-$1$ quaternion-domain GEK matrix, which incorporates all pairwise distances and ADoA information among the nodes, is given in (32), shown at the top of the next page. Assuming that measured values of all pairwise distance and ADoA parameters appearing in (31) are available, the quaternion-domain GEK matrix with measurement errors, denoted as $\tilde{\bm{K}}_{\mathrm{q}}$, can be constructed. Accordingly, the estimate of the quaternion edge vector $\bm{\nu}$ is given as
@@
The SVD of $\tilde{\bm{K}}_{\mathrm{q}}$ is computed using the QSVD method introduced in Section II-B. Since $\tilde{\bm{K}}_{\mathrm{q}}$ has rank $1$ by construction, its equivalent complex representation has rank $2$ with identical eigenvalues [33]. Furthermore, because the quaternion-domain GEK $\tilde{\bm{K}}_{\mathrm{q}}$ is regular in our setting, the quaternion eigendecomposition is effectively equivalent to the QSVD.
@@
The computational complexity of the QD-SMDS algorithm is dominated by the QSVD operation, which corresponds to a truncated SVD of the complex-valued representation of the quaternion kernel matrix in (32), having size $2M\times 2M$. Therefore, the asymptotic complexity is $O\left(2(2M)^{2}\right)=O\left(M^{2}\right)=O\left(N_{A}^{2}(N_{A}+N_{\mathrm{T}})^{2}\right)$, which is the same as that of the SMDS algorithm.